If a sphere has radius 5 m. what is the surface area of the sphere?
Answer by jim_thompson5910(35256)

SA = 4*pi*r^2

SA = 4*pi*5^2

SA = 4*pi*25

SA = 4*25*pi

SA = 100pi <---- Exact Surface Area

SA = 100*(3.14) ... using 3.14 for pi

SA = 314 <---- Approximate Surface Area

Note: units for the answer are square meters.